# A Survey on Wi-Fi Sensing Generalizability: Taxonomy, Techniques, Datasets, and Future Research Prospects

### v2

@@ IV-A Data, Data, Data
Although researchers have made significant efforts to close this gap by collecting and releasing datasets like XRF55 [70] and Widar 3.0 [67], these datasets still suffer from constraints in terms of collection scenarios, action categories, and participant diversity. When compared to the scale of video, image, or text datasets, Wi-Fi sensing datasets lag behind by several orders of magnitude.
Collecting and annotating Wi-Fi sensing data is inherently time-consuming and labor-intensive. To address this bottleneck, synthetic data generation emerges as a promising solution. By producing large volumes of realistic data, synthetic methods can support the training of models with substantial parameters, potentially revolutionizing Wi-Fi sensing research. However, several open challenges remain to be addressed:
@@
    How to synthesize diverse human body models? Is it more effective to use game engines like Unity or Unreal Engine to create diverse human models? Alternatively, could tools such as SMPL [308] or MANO [309] be utilized to extract human body models from large-scale video datasets for synthetic data generation?
@@
    How to bridge the gap between synthetic and real data? Could Generative Adversarial Networks (GANs) [252] help enhance the realism of synthetic features or adapt synthetic data to better match real-world characteristics? What role could domain adaptation and domain-invariant feature learning play in aligning synthetic and real data to improve model performance?
Addressing these challenges offers a pathway to creating realistic, diverse, and scalable synthetic Wi-Fi sensing datasets. By bridging the data gap, the Wi-Fi sensing field could achieve progress comparable to the rapid advancements seen in video, image, and text recognition, ultimately unlocking new frontiers for this promising domain.
@@ IV-B Wi-Fi Sensing Foundation Model Pre-Training
    Wi-Fi sensing is significantly influenced by environmental factors and device configurations. Incorporating such information could enable the foundation model to reason more effectively about environmental changes and adapt to new scenarios. Should this integration be achieved through explicit embeddings of environmental and device parameters, or should the model be designed to infer these characteristics implicitly from the data?
Addressing these questions will be critical for developing a robust and generalizable Wi-Fi sensing foundation model that can serve as a backbone for various downstream applications.
@@ IV-C When Wi-Fi Sensing Meets Large Multimodal Models
    How can Wi-Fi data be aligned with the modalities handled by large multimodal models? Wi-Fi sensing data typically exists in time-series or signal-based formats, which differ significantly from image or text data commonly processed by multimodal models. Should the raw Wi-Fi data be transformed into spectrograms or other image-like representations for easier integration, or is it better to retain the time-series format and develop alignment techniques specific to such data? Furthermore, what pre-processing or feature extraction methods are necessary to achieve optimal performance when aligning Wi-Fi data with existing modalities?
@@
    What role does multimodal fusion play in Wi-Fi sensing? Large multimodal models excel at integrating and reasoning over diverse data sources. In Wi-Fi sensing, how can complementary modalities such as video, audio, or inertial measurements (e.g., IMU data) be effectively fused with Wi-Fi data? What techniques can ensure seamless fusion while addressing challenges like modality-specific noise, synchronization issues, and computational overhead?
Addressing these challenges will advance the integration of Wi-Fi sensing with large multimodal models and push the boundaries of their application across various tasks and environments. This direction holds great potential for accelerating progress in Wi-Fi sensing by leveraging the capabilities of state-of-the-art AI models.
IV-D System Deployment and Service Models
As Wi-Fi sensing technologies advance and large multimodal models are integrated into Wi-Fi sensing applications, an essential next step involves addressing deployment and service models for these systems. Several critical questions arise when considering how to effectively and efficiently deploy such systems in real-world settings:
@@ IV-D System Deployment and Service Models
    Large multimodal models often require high computational power and storage. Compression techniques like pruning [319], quantization [320], or knowledge distillation [321] can help balance performance and efficiency. How can these methods be tailored for different hardware (e.g., edge devices or mobile phones) to ensure speed without sacrificing too much performance?
@@
    How to manage network bandwidth, data traffic, and latency during operation? Network constraints, particularly with real-time data transmission, pose challenges. How can continuous data flow between edge devices and servers be optimized? Strategies like edge computing [322], which processes data locally, could reduce latency and bandwidth usage but introduce complexity in resource management and synchronization.
@@
    Should model updates be centralized or on-demand at the edge? Model updates can either be pushed centrally from a server or performed on-demand at the edge based on local needs. Centralized updates ensure consistency but may face bandwidth and computation constraints. Edge updates offer quick adaptability but can complicate coordination, version control, and synchronization.
@@
    What service and pricing strategies should be adopted? Deciding on the right service and pricing models is vital for commercialization. Should services be cloud-based, subscription-based, or pay-per-use? How can pricing be adapted based on usage, such as the number of devices, frequency of access, or task complexity? Should tiered services be offered, with basic models for simpler tasks and premium ones for more complex applications?
Addressing these questions will help define how Wi-Fi sensing services can be deployed and scaled efficiently in real-world applications, ultimately determining the practicality and success of Wi-Fi sensing systems across diverse industries and environments.
@@ V Conclusion
In this survey, we systematically reviewed and categorized Wi-Fi sensing generalization studies published between 2014 and 2025, following the end-to-end pipeline of Wi-Fi sensing systems—spanning device deployment, model construction, and model deployment. We summarized key techniques proposed in each stage to address the challenges of domain variability. In addition, we provided a comprehensive overview of existing publicly available datasets, highlighting their domain diversity and applicability to generalization research.
@@
[…]

### v4

@@ IV-A Data, Data, Data
Recent initiatives are actively narrowing the data gap; for instance, CSI-Bench [323] achieves a milestone in realism with 461 hours of recordings, XRF55 [208] significantly expands the sample count to 42,900, and OctoNet [264] comprises 12 heterogeneous modalities. Despite these advancements in data duration and environmental complexity, Wi-Fi sensing datasets still lack the vast scale of CV benchmarks like ImageNet or Kinetics. The gap remains evident not only in raw sample counts but also in the breadth of action categories and participant demographics, emphasizing the need for continued expansion in data diversity.
Collecting and annotating Wi-Fi sensing data is inherently time-consuming and labor-intensive, creating a data-scarcity bottleneck for deep learning. To address this, synthetic data generation has emerged as a promising solution. By leveraging high-fidelity 3D modeling and computational electromagnetics, researchers can simulate diverse human-environment interactions. For example, by placing dynamic 3D human models into various virtual environments and employing ray-tracing algorithms to emulate signal propagation, realistic Wi-Fi CSI data can be synthesized. This approach provides the massive datasets required to train large-scale foundation models, potentially revolutionizing the scalability of Wi-Fi sensing. However, several open challenges, such as the domain gap between simulated and physical signals, remain to be addressed.
@@
    How to synthesize diverse human body models? Is it more effective to use game engines like Unity or Unreal Engine to create diverse human models? Alternatively, could tools such as SMPL [140] or MANO [173] be utilized to extract human body/hand models from large-scale video datasets for synthetic data generation?
@@
    How to bridge the gap between synthetic and real data? Could Generative Adversarial Networks [63] help enhance the realism of synthetic features or adapt synthetic data to better match real-world characteristics? What role could domain adaptation and domain-invariant feature learning play in aligning synthetic and real data to improve model performance?
Addressing these challenges offers a pathway to creating realistic, diverse, and scalable synthetic Wi-Fi sensing datasets. By bridging the data gap, the Wi-Fi sensing field could achieve progress comparable to the rapid advancements seen in video, image, and text recognition, ultimately unlocking new frontiers for this promising domain. We consider a strategic roadmap for future research in synthetic data generation:
  • Short-term Goals (1–2 years): The primary focus should be on enhancing signal realism. Leveraging advanced generative models, such as Diffusion Models or Generative Adversarial Networks, researchers can simulate more authentic multi-path effects and hardware-induced noise. The prioritized challenge lies in establishing a precise “signal-to-physical-entity” mapping, ensuring that synthesized CSI samples are indistinguishable from real-world data in both the frequency and time domains.
  • Long-term Goals (3–5 years): The ultimate objective is to realize large-scale, multi-environment closed-loop simulation systems. This involves developing sophisticated simulation engines capable of automatically generating complex 3D indoor scenes and simulating dynamic human-environment interactions. This leap from “single-sample synthesis” to “full-scene simulation” will be critical. The core challenge remains overcoming the “Sim-to-Real” domain gap, ensuring that models trained entirely on simulated data can be seamlessly deployed in physical hardware environments.
@@ IV-B Wi-Fi Sensing Foundation Model Pre-Training
    Wi-Fi sensing is significantly influenced by environmental factors, human body factors, and device configurations. Incorporating such information could enable the foundation model to reason more effectively about environmental changes and adapt to new scenarios. Should this integration be achieved through explicit embeddings of environmental parameters, human body factors, and device configurations, or should the model be designed to infer these characteristics implicitly from the data?
Addressing these questions will be critical for developing a robust and generalizable Wi-Fi sensing foundation model that can serve as a backbone for various downstream applications. For the pre-training of large-scale Wi-Fi sensing models, we outline the following roadmap:
  • Short-term Goals (1–2 years): The primary objective is to construct task-agnostic universal feature extractors. By leveraging existing medium-scale datasets, researchers can employ self-supervised learning strategies, such as Masked Signal Modeling, to learn fundamental feature representations that are robust against device heterogeneity.
  • Long-term Goals (3–5 years): The ultimate aim is to train perception foundation models with billions of parameters. By integrating ultra-large-scale unlabeled real-world data with synthetic datasets, the model will achieve powerful zero-shot generalization capabilities. The key challenges involve addressing the massive computational costs associated with processing large-scale CSI data, as well as coordinating globally distributed devices for pre-training, while strictly adhering to user privacy protocols.
@@ IV-C When Wi-Fi Sensing Meets Large Multimodal Models
    How can Wi-Fi data be aligned with the modalities handled by large multimodal models (LMMs)? Wi-Fi sensing data typically exists in time-series formats, which differ significantly from the visual or textual tokens processed by LMMs. Building on proven strategies such as converting CSI into spectrograms or image-like representations [275, 99], and inspired by cross-modal translation works like WiFi2Radar [153], it is evident that aligning Wi-Fi with more ‘visual’ modalities is a viable path for leveraging LMMs. However, while these transformation techniques have shown success in specific tasks, they have yet to become a mainstream standard for universal LMM integration. This raises a critical research question: should we continue to adapt Wi-Fi data to fit visual architectures, or is it more effective to retain its native time-series format and develop specific alignment techniques (e.g., cross-modal contrastive learning) tailored for LMMs? Furthermore, identifying which pre-processing or feature extraction methods are essential to bridge the semantic gap between Wi-Fi signals and existing modalities remains a key bottleneck for achieving optimal performance in multimodal sensing.
@@
    What role does multimodal fusion play in Wi-Fi sensing? In scenarios where a multimodal sensing setup is available, such as Wi-Fi transceivers co-deployed with cameras or IMUs, multimodal fusion plays a vital role in enhancing performance, robustness, and sensing coverage. Large multimodal models excel at integrating these diverse data sources by aligning them in a shared semantic space together with techniques like ALBEF [118], Cross-Attention Mechanism [28], and Mixture of Experts [265]. However, achieving seamless fusion requires advanced techniques to address modality-specific noise, temporal synchronization, and the high computational overhead associated with processing multi-source data.
Addressing these challenges will advance the integration of Wi-Fi sensing with large multimodal models and push the boundaries of their application across various tasks and environments. This direction holds great potential for accelerating progress in Wi-Fi sensing by leveraging the capabilities of state-of-the-art AI models. To facilitate the integration of Wi-Fi sensing with LMMs, we envision the following roadmap:
  • Short-term Goals (1–2 years): The immediate priority is to establish efficient semantic alignment mechanisms between Wi-Fi signals and other modalities like vision or text. Drawing inspiration from cross-modal translation frameworks such as WiFi2Radar, Wi-Fi signals can be transformed into LMM-compatible tokens or spectrograms. Research should prioritize resolving the spatio-temporal misalignment between Wi-Fi time-series data and visual imagery.
  • Long-term Goals (3–5 years): The long-term objective is to develop native sensing agents with inherent wireless perception capabilities. These agents will enable LMMs to perform direct logical reasoning based on Wi-Fi signals, facilitating complex, real-world tasks. The fundamental challenge lies in bridging the semantic gap between the low-level physical properties of wireless signals and the high-level reasoning logic of large models.
IV-D System Deployment
As Wi-Fi sensing technologies advance and large multimodal models are integrated into Wi-Fi sensing applications, an essential next step involves addressing system deployment. Typically, this deployment could follow a hierarchical ‘Cloud-to-Edge’ pipeline: initially, a Wi-Fi sensing model is pre-trained on extensive data in the cloud to learn generalizable representations. This model is then compressed and transmitted to resource-constrained edge devices (e.g., routers), where it performs lightweight on-device adaptation to align with the unique multipath profiles and user behaviors of the target environment. However, several critical questions arise when considering how to effectively and efficiently deploy such systems in real-world settings:
@@ IV-D System Deployment
    Efficient Inference: Deploying large models on COTS chips (e.g., ESP32) requires handling Wi-Fi’s frequency-domain sparsity. The core research question is: How to design compression techniques such as structure-aware pruning [71], quantization [93], and knowledge distillation [80] that identify and retain only those subcarriers sensitive to human motion, thereby minimizing computational overhead while preserving sensing granularity?
@@
    On-device training and fine-tuning. Wi-Fi sensing is uniquely susceptible to environmental shifts; even minor furniture changes alter multipath profiles. Centralized updates often fail to address these site-specific fingerprints. Therefore, a more effective strategy is environment-aware edge adaptation, deploying a foundation model centrally while performing on-demand fine-tuning with techniques like on-device training [130]. This allows rapid calibration to local environments and user patterns without massive data backhaul.
@@
    Sensing-Communication Coexistence. Continuous CSI streaming significantly consumes bandwidth, affecting concurrent communication tasks. The critical challenge is optimizing resource allocation. Specifically, during network congestion, how can the system adaptively downsample CSI rates or perform temporal-spectral compression to maintain real-time sensing (e.g., fall detection) without degrading communication Quality of Service? Future strategies must focus on ISAC-aware scheduling, ensuring sensing pilots and data packets are synchronized based on task urgency.
@@
    Adversarial Resilience. Sensing models are vulnerable to physical adversarial attacks that manipulate system output [134]. How can we protect systems against channel interference and signal alteration? What approaches, such as robust adversarial training or error correction protocols, can be adopted to mitigate such risks? Additionally, recent advances in Generative AI may offer new solutions for physical-layer security, as seen in DFSS [213].
  5. 5.
    In consumer scenarios, requiring users to manually collect and annotate data is often infeasible. Practical deployment therefore depends on minimizing this burden through two main strategies. First, implicit feedback loops can leverage user corrections, e.g., manually switching a light after a false detection, as supervision signals. Second, cross-modal supervision can use auxiliary sensors, such as smartphone IMUs, to automatically generate labels during the initial adaptation phase, eliminating the need for manual annotation. Despite these strategies, significant challenges remain: feedback- or sensor-derived labels may be noisy or incomplete, privacy concerns can restrict the use of auxiliary sensing modalities, and long-term deployments must contend with concept drift caused by changes in user behavior or the environment. These issues underscore the need for labeling-efficient, privacy-preserving, and user-friendly data collection mechanisms, which continue to be an open research direction for deployable Wi-Fi sensing systems.
Addressing these questions will help define how Wi-Fi sensing services can be deployed and scaled efficiently in real-world applications, ultimately determining the practicality and success of Wi-Fi sensing systems across diverse industries and environments. Regarding the practical deployment and long-term evolution of Wi-Fi sensing, we outline the following roadmap:
  • Short-term Goals (1–2 years): The focus is on establishing lightweight adaptive deployment mechanisms and standardized sensing interfaces. Research should prioritize low-power, calibration-free adaptive algorithms that enable systems to rapidly adjust to new environments via weak supervision signals (e.g., user-driven corrections). Simultaneously, promoting the early adoption of standards such as IEEE 802.11bf is crucial to unifying CSI extraction interfaces across hardware vendors, thereby lowering the deployment barriers caused by device heterogeneity.
  • Long-term Goals (3–5 years): The ultimate objective is to realize a large-scale, privacy-preserving “Sensing-as-a-Service” ecosystem. This involves integrating Wi-Fi sensing into the fundamental infrastructure of smart cities.
@@ V Conclusion
In this survey, we systematically reviewed and categorized Wi-Fi sensing generalization studies published between 2015 and 2025, following the end-to-end pipeline of Wi-Fi sensing systems—spanning device deployment, signal preprocessing, feature learning, and model deployment. We summarized key techniques proposed in each stage to address the challenges of domain variability. In addition, we provided a comprehensive overview of existing publicly available datasets, highlighting their domain diversity and applicability to generalization research.
@@
To further promote the development of Wi-Fi sensing, we have initiated the Sensing Dataset Platform (SDP: http://www.sdp8.org ), an open dataset and model sharing hub designed to foster collaboration between academia and industry, enabling researchers worldwide to contribute and access high-quality resources. Given the rapid pace in this field, we will continuously maintain and update relevant resources at https://github.com/aiotgroup/awesome-wireless-sensing-generalization.